OpenStudy (anonymous):

simplify: 5^5log(base1/5)x

OpenStudy (anonymous):

first simplify log(1/5)(x) log(1/5)(x) = log(x)/log(1/5) =log(x)/(-log(5)) =-log(x)/log(5) = log(1/x)/log(5) =log(base5)(1/x) so 5^5log(1/5)(x) = 5^5log5(1/x) = 5^log5(1/x^5) =1/x^5 (because e^(loge(a))=a

OpenStudy (anonymous):

1/5 is the BASE. You cannot use the above mentioned property!

OpenStudy (anonymous):

It is not log1/5*x , it is \[\log_{1/5} x\]

OpenStudy (anonymous):

if you dont know, learn it. the formula is log(base a)(b) = log(b)/log(a) (for any fictious base)

OpenStudy (anonymous):

\[\log _{b}^{a} = \log(a)/\log(b)\]

OpenStudy (anonymous):

penpal is correct..that is how the problem is written